# A Debiased Estimator for the Mediation Functional in Ultra-High-Dimensional Setting in the Presence of Interaction Effects

### v2

@@ 4.1 Bias Analysis
In this section, we present the bias structure of the debiased estimator $\hat{\theta}_{0}$ and show that the bias is asymptotically negligible, i.e., it decays at a faster rate than $n^{-1/2}$. We would like to point out that debiasing a high-dimensional parameter does not make it unbiased, i.e., it does not remove the bias completely, rather reduces the bias enough to achieve $\sqrt{n}$-consistency (e.g., see (Zhang and Zhang, , 2014; Van de Geer et al., , 2014; Javanmard and Montanari, , 2014)). To analyze the remaining bias of $\hat{\theta}_{0}$, first, recall the definitions $b:=\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}$, $\hat{b}:=\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}$, $\hat{\phi}_{1}:=(\hat{\beta}_{1}^{\top},\hat{\gamma}_{1}^{\top})^{\top}$, and the contrast $a:=(\bar{X}^{\top}\quad(\hat{\mathbf{B}}_{0}\bar{X})^{\top})^{\top}$ from the previous section. The following proposition establishes an upper bound on the bias of $\hat{\theta}_{0}$.
@@
$|\Delta_{n}|\leq\|\mathbf{B}_{0}\bar{X}-\hat{\mathbf{B}}_{0}\bar{X}\|_{\infty}\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}+\|a-\mathbf{W}_{t}^{\top}\tau_{1}\|_{\infty}\|\hat{\phi}_{1}-\phi_{1}\|_{1}+\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\|_{\infty}\|\hat{b}-b\|_{1}\,.$ (4.1)
See Appendix A.1 for a proof. Observe that each of the three terms in the upper bound of $|\Delta_{n}|$ has a product form, i.e., the product of an estimation and/or an approximation error. More specifically, the first summand is the product of the estimation error of $\mathbf{B}_{0}$ (along the direction of $\bar{X}$) and $\gamma_{1}$, the second summand is the product of the estimation error of $\phi_{1}$ and the approximation error of $a$ by $\mathbf{W}_{t}^{\top}\tau_{1}$, and the last summand is the product of the estimation error of $b$ and the approximation error of $\bar{X}$ by $\mathbf{X}_{c}^{\top}\tau_{2}$. Whilst each of these individual estimation/approximation errors can be significantly larger than $n^{-1/2}$ in order, the product form is what makes each summand asymptotically negligible. In Section 5, we provide sufficient conditions on the sparsity of parameters and the distribution of the error terms, under which $|\Delta_{n}|=o_{p}(n^{-1/2})$, and $\hat{\theta}_{0}$ is $\sqrt{n}$-consistent and asymptotically normal.
@@ 4.2 Connection to Fixed-Dimensional Setting
In a fixed dimensional setting (i.e., when the number of covariates as well as potential mediators are fixed), Tchetgen Tchetgen and Shpitser, (2012) provided a debiased estimator of the mediation functional which was constructed based on the influence function of the parameter of interest. Our proposed estimator can be considered as the counterpart of their approach in the UHD setting. The debiased estimator, as proposed in (Tchetgen Tchetgen and Shpitser, , 2012), can be written as
$\textstyle\hat{\theta}^{IF}$ $=\mathbb{E}_{n}\bigg{[}\hat{\mu}_{10}(X)+\frac{A}{\hat{p}(A=1\mid X)}\frac{\hat{p}(M\mid A=0,X)}{\hat{p}(M\mid A=1,X)}\Big{(}Y-\hat{\mu}_{1}(X,M)\Big{)}+\frac{1-A}{\hat{p}(A=0\mid X)}\Big{(}\hat{\mu}_{1}(X,M)-\hat{\mu}_{10}(X)\Big{)}\bigg{]}\,,$ (4.2)
@@
We now show that our estimator $\hat{\theta}_{0}$, obtained via Algorithm 1, is structurally similar to $\hat{\theta}^{IF}$ when the underlying data generating process is linear. Consider the following specific nuisance function choices $\hat{\mu}_{1}(X,M)=X^{\top}\hat{\beta}_{1}+M^{\top}\hat{\gamma}_{1}$, and $\hat{\mu}_{10}(X)=X^{\top}(\hat{\beta}_{1}+\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}})$. Then, estimator $\hat{\theta}^{IF}$ will be of the form of
$\hat{\theta}^{IF}$ $=\mathbb{E}_{n}\bigg{[}X^{\top}\left(\hat{\beta}_{1}+\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right)+\frac{A}{\hat{\pi}(X)}\frac{\hat{p}(M\mid A=0,X)}{\hat{p}(M\mid A=1,X)}\left(Y-X^{\top}\hat{\beta}_{1}-M^{\top}\hat{\gamma}_{1}\right)$
$\quad+\frac{1-A}{1-\hat{\pi}(X)}\left(X^{\top}\hat{\beta}_{1}+M^{\top}\hat{\gamma}_{1}-X^{\top}\left(\hat{\beta}_{1}+\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right)\right)\bigg{]}$
$=\mathbb{E}_{n}\left[X^{\top}\left(\hat{\beta}_{1}+\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right)+\omega_{1}A\left(Y-X^{\top}\hat{\beta}_{1}-M^{\top}\hat{\gamma}_{1}\right)+\omega_{2}(1-A)\left(M^{\top}\hat{\gamma}_{1}-X^{\top}\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right)\right]$
$=\bar{X}^{\top}\left(\hat{\beta}_{1}+\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right)+\frac{1}{n}\sum_{\{i:A_{i}=1\}}\omega_{1,i}\left(Y_{i}-X_{i}^{\top}\hat{\beta}_{1}-M_{i}^{\top}\hat{\gamma}_{1}\right)+\frac{1}{n}\sum_{\{i:A_{i}=0\}}\omega_{2,i}\left(M_{i}^{\top}\hat{\gamma}_{1}-X_{i}^{\top}\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right)\,,$ (4.3)
@@
$\hat{\pi}(X_{i})=\hat{p}(A=1\mid X_{i}),\ \ \omega_{1,i}=\frac{1}{\hat{\pi}(X_{i})}\frac{\hat{p}(M_{i}\mid A_{i}=0,X_{i})}{\hat{p}(M_{i}\mid A_{i}=1,X_{i})},\quad\omega_{2,i}=\frac{1}{1-\hat{\pi}(X_{i})}.$
@@
$\hat{\theta}_{0}=\bar{X}^{\top}(\hat{\beta}_{1}+\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}})+\sum_{\{i:A_{i}=1\}}\tau_{1,i}\left(Y_{i}-X_{i}^{\top}\hat{\beta}_{1}-M_{i}^{\top}\hat{\gamma}_{1}\right)+\sum_{\{i:A_{i}=0\}}\tau_{2,i}\left(M_{i}^{\top}\hat{\gamma}_{1}-X_{i}^{\top}\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right)\,$ (4.4)
@@
We next consider the following reparametrization of the estimator of the mediation functional in Equation (4.2), as recently studied in (Ghassami, , 2024):
$\hat{\theta}^{mr}=\mathbb{E}_{n}\left[\hat{\mu}_{10}(X)+\frac{(1-A)}{\hat{\varpi}_{1}(X)}\left\{\hat{\mu}_{1}(M,X)-\hat{\mu}_{10}(X)\right\}+\frac{A\hat{\varpi}_{2}(M,X)}{1-\hat{\varpi}_{1}(X)}\left\{Y-\hat{\mu}_{1}(M,X)\right\}\right],$ (4.5)
where,
$\varpi_{1}(X)$ $=p(A=0\mid X),\quad\varpi_{2}(M,X)=\frac{p(M\mid A=0,X)}{p(M\mid A=1,X)}.$
It was shown in (Ghassami, , 2024) that under this reparametrization in the fixed-dimensional setting, the bias has the following form.
$\mathbb{E}\left[\hat{\theta}^{mr}\right]-\theta_{0}=$ $\mathbb{E}\left[\left\{\frac{(1-A)}{\hat{\varpi}_{1}(X)}-1\right\}\left\{\mu_{10}(X)-\hat{\mu}_{10}(X)\right\}\right]$
$\qquad+\mathbb{E}\left[\left\{\frac{A\hat{\varpi}_{2}(M,X)}{1-\hat{\varpi}_{1}(X)}-\frac{(1-A)}{\hat{\varpi}_{1}(X)}\right\}\left\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\right\}\right].$ (4.6)
In the fixed-dimension, by the Hölder’s inequality and plugging in $\hat{\mu}_{1}(X,M)=X^{\top}\hat{\beta}_{1}+M^{\top}\hat{\gamma}_{1}$, and $\hat{\mu}_{10}(X)=X^{\top}(\hat{\beta}_{1}+\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}})$, Equation (4.6) can be upper bounded by
@@
$\leq\mathbb{E}\left[\left\{\left\|X^{\top}\left(\frac{A\hat{\varpi}_{2}(M,X)}{1-\hat{\varpi}_{1}(X)}-\frac{1-A}{\hat{\varpi}_{1}(X)}\right)\right\|_{\infty}+\left\|X^{\top}\left(\frac{1-A}{\hat{\varpi}_{1}(X)}-1\right)\right\|_{\infty}\right\}\|\hat{\beta}_{1}-\beta_{1}\|_{1}\right]$
$\quad+\mathbb{E}\left[\left\|M^{\top}\left(\frac{A\hat{\varpi}_{2}(M,X)}{1-\hat{\varpi}_{1}(X)}-\frac{(1-A)}{\hat{\varpi}_{1}(X)}\right)\right\|_{\infty}\left\|\gamma_{1}-\hat{\gamma}_{1}\right\|_{1}\right]$
$\quad+\mathbb{E}\left[\left\|X^{\top}\left(\frac{1-A}{\hat{\varpi}_{1}(X)}-1\right)\right\|_{\infty}\left\|\mathbf{B}_{0}^{\top}\gamma_{1}-\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right\|_{1}\right].$ (4.7)
@@
$|\Delta_{n}|$ $\leq\|a-\mathbf{W}^{\top}\tau_{1}\|_{\infty}\|\hat{\phi}_{1}-\phi_{1}\|_{1}+\|(\hat{\mathbf{B}}_{0}-\mathbf{B}_{0})\bar{X}\|_{\infty}\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}+\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\|_{\infty}\|\hat{b}-b\|_{1}$
$=\|a-\mathbf{W}^{\top}\tau_{1}\|_{\infty}\|\hat{\beta}_{1}-\beta_{1}\|_{1}$
$\quad+\left(\|a-\mathbf{W}^{\top}\tau_{1}\|_{\infty}+\|(\hat{\mathbf{B}}_{0}-\mathbf{B}_{0})\bar{X}\|_{\infty}\right)\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$
@@
It is apparent that the bias structures of $\hat{\theta}^{mr}$ and $\hat{\theta}_{0}$ are similar, as both of them are the sum of three terms, and each of them is the product of two errors. We summarize the similarity in the following table.
$\begin{array}[]{c c c}\text{High-dimensional setting}&&\text{Fixed-dimensionalsetting}\\ \hline\cr\|a-\mathbf{W}^{\top}\tau_{1}\|_{\infty}\|\hat{\beta}_{1}-\beta_{1}\|_{1}&\longleftrightarrow&\begin{matrix}\big{(}\|X^{\top}\left(\frac{A\hat{\varpi}_{2}(M,X)}{1-\hat{\varpi}_{1}(X)}-\frac{1-A}{\hat{\varpi}_{1}(X)}\right)\|_{\infty}\\ ~{}~{}~{}~{}~{}~{}+\|X^{\top}\left(\frac{1-A}{\hat{\varpi}_{1}(X)}-1\right)\|_{\infty}\big{)}\|\beta_{1}-\hat{\beta}_{1}\|_{1}\end{matrix}\\ \left(\|a-\mathbf{W}^{\top}\tau_{1}\|_{\infty}+\|(\hat{\mathbf{B}}_{0}-\mathbf{B}_{0})\bar{X}\|_{\infty}\right)\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}&\longleftrightarrow&\|M^{\top}\left(\frac{A\hat{\varpi}_{2}(M,X)}{1-\hat{\varpi}_{1}(X)}-\frac{(1-A)}{\hat{\varpi}_{1}(X)}\right)\|_{\infty}\left\|\gamma_{1}-\hat{\gamma}_{1}\right\|_{1}\\ \|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\|_{\infty}\|\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}-\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}\|_{1}&\longleftrightarrow&\|X^{\top}\left(\frac{1-A}{\hat{\varpi}_{1}(X)}-1\right)\|_{\infty}\|\mathbf{B}_{0}^{\top}\gamma_{1}-\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\|_{1}\\ \end{array}$
Relation to balancing: We also draw a connection between our debiasing technique and the balancing techniques in the literature of causal inference. In the influence function-based estimator for ATE, the correction term (the term added to the initial estimator to reduce bias) contains the propensity score as a nuisance function. Imai and Ratkovic, (2014) showed that the propensity score satisfies a certain balancing property, based on which they proposed a balancing estimator for the propensity score. Zubizarreta, (2015) proposed an optimization-based approach for achieving the said balance. In the high-dimensional setting, Athey et al., (2018) illustrated that the standard debiasing technique of penalized regression-based estimator (e.g., the method proposed earlier in (Javanmard and Montanari, , 2014; Van de Geer et al., , 2014; Zhang and Zhang, , 2014)) is inherently related to the balancing approach of Zubizarreta, (2015). However, none of these approaches is directly applicable in the presence of mediators for assessing the mediation functional. Recently, Ghassami, (2024) presented a counterpart of the balancing property and balancing estimator of Imai and Ratkovic, (2014) for mediation functional in a fixed-dimensional setting. A version of the balancing property in that work can be achieved by minimizing $\left\|\frac{1-A}{\hat{\varpi}_{1}(X)}-1\right\|_{\infty}$ and $\left\|\frac{A\hat{\varpi}_{2}(M,X)}{1-\hat{\varpi}_{1}(X)}-\frac{(1-A)}{\hat{\varpi}_{1}(X)}\right\|_{\infty}$. As described above, the counter part of these two quantities in our setting are $\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\|_{\infty}$ and $\|a-\mathbf{W}_{t}^{\top}\tau_{1}\|_{\infty}$, respectively. Hence, the two optimizations (3.1) and (3.2) performed in our proposed estimation strategy can be viewed as adaptations of the balancing of Ghassami, (2024) in the UHD setting.

### v3

@@ 4.1 Bias Analysis
In this section, we present the bias structure of the debiased estimator $\hat{\theta}_{0}$ and show that the bias is asymptotically negligible, i.e., it decays at a faster rate than $n^{-1/2}$. We would like to point out that debiasing a high-dimensional parameter does not make it unbiased, i.e., it does not remove the bias completely, but rather reduces the bias enough to achieve $\sqrt{n}$-consistency (e.g., see (Zhang and Zhang, , 2014; Van de Geer et al., , 2014; Javanmard and Montanari, , 2014)). To analyze the remaining bias of $\hat{\theta}_{0}$, first, recall the definitions $b:=\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}$, $\hat{b}:=\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}$, $\hat{\phi}_{1}:=(\hat{\beta}_{1}^{\top},\hat{\gamma}_{1}^{\top})^{\top}$, and the contrast $a:=(\bar{X}^{\top}\quad(\hat{\mathbf{B}}_{0}\bar{X})^{\top})^{\top}$ from the previous section. The following proposition establishes an upper bound on the bias of $\hat{\theta}_{0}$.
@@
$|\Delta_{n}|$ $\leq\|\mathbf{B}_{0}\bar{X}-\hat{\mathbf{B}}_{0}\bar{X}\|_{\infty}\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$ (4.1)
$~~~~~~~~~~~~~+\|a-\mathbf{W}_{t}^{\top}\tau_{1}\|_{\infty}\|\hat{\phi}_{1}-\phi_{1}\|_{1}+\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\|_{\infty}\|\hat{b}-b\|_{1}\,.$
See Appendix A.1 for a proof. Observe that each of the three terms in the upper bound of $|\Delta_{n}|$ has a product form, i.e., the product of an estimation and/or an approximation error. More specifically, the first summand is the product of the estimation error of $\mathbf{B}_{0}$ (along the direction of $\bar{X}$) and $\gamma_{1}$, the second summand is the product of the estimation error of $\phi_{1}$ and the approximation error of $a$ by $\mathbf{W}_{t}^{\top}\tau_{1}$, and the last summand is the product of the estimation error of $b$ and the approximation error of $\bar{X}$ by $\mathbf{X}_{c}^{\top}\tau_{2}$. Whilst each of these individual estimation/approximation errors can be significantly larger than $n^{-1/2}$ in order, the product form is what makes each summand asymptotically negligible. In Section 5, we provide sufficient conditions on the sparsity of parameters and the distribution of the error terms, under which $|\Delta_{n}|=o_{p}(n^{-1/2})$, and $\hat{\theta}_{0}$ is $\sqrt{n}$-consistent and asymptotically normal, as $\sqrt{n}V_{n}$ will be shown to converge to standard Gaussian distribution.
@@ 4.2 Connection to Fixed-Dimensional Setting
In a fixed dimensional setting (i.e., when the number of covariates as well as potential mediators is fixed), Tchetgen Tchetgen and Shpitser, (2012) provided a debiased estimator of the mediation functional, which was constructed based on the influence function of the parameter of interest. Our proposed estimator can be considered as the counterpart of their approach in the UHD setting. The debiased estimator, as proposed by Tchetgen Tchetgen and Shpitser, (2012), can be written as
$\textstyle\hat{\theta}^{IF}$ $=\mathbb{E}_{n}\bigg{[}\hat{\mu}_{10}(X)+\frac{A}{\hat{p}(A=1\mid X)}\frac{\hat{p}(M\mid A=0,X)}{\hat{p}(M\mid A=1,X)}\Big{(}Y-\hat{\mu}_{1}(X,M)\Big{)}$ (4.2)
$~~~~~~~~~~~~~~~~~~~~~~~~~~+\frac{1-A}{1-\hat{p}(A=1\mid X)}\Big{(}\hat{\mu}_{1}(X,M)-\hat{\mu}_{10}(X)\Big{)}\bigg{]}\,$,
where $\hat{\mu}_{10}(X)=\int\hat{\mu}_{1}(X,M)\hat{p}(M\mid A=0,X)dM$.
@@
$\hat{\theta}^{IF}$ $=\mathbb{E}_{n}\bigg{[}X^{\top}\left(\hat{\beta}_{1}+\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right)+\frac{A}{\hat{p}(A=1\mid X)}\frac{\hat{p}(M\mid A=0,X)}{\hat{p}(M\mid A=1,X)}\left(Y-X^{\top}\hat{\beta}_{1}-M^{\top}\hat{\gamma}_{1}\right)$ (4.3)
[…]
$=\mathbb{E}_{n}\bigg{[}X^{\top}\left(\hat{\beta}_{1}+\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right)+\omega_{1}A\left(Y-X^{\top}\hat{\beta}_{1}-M^{\top}\hat{\gamma}_{1}\right)$
[…]
$=\bar{X}^{\top}\left(\hat{\beta}_{1}+\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right)+\frac{1}{n}\sum_{\{i:A_{i}=1\}}\omega_{1,i}\left(Y_{i}-X_{i}^{\top}\hat{\beta}_{1}-M_{i}^{\top}\hat{\gamma}_{1}\right)$
[…]
@@
$\omega_{1,i}=\frac{1}{\hat{p}(A_{i}=1\mid X_{i})}\frac{\hat{p}(M_{i}\mid A_{i}=0,X_{i})}{\hat{p}(M_{i}\mid A_{i}=1,X_{i})},\quad\omega_{2,i}=\frac{1}{1-\hat{p}(A_{i}=1\mid X_{i})}.$
@@
$\hat{\theta}_{0}$ $=\bar{X}^{\top}(\hat{\beta}_{1}+\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}})+\sum_{\{i:A_{i}=1\}}\tau_{1,i}\left(Y_{i}-X_{i}^{\top}\hat{\beta}_{1}-M_{i}^{\top}\hat{\gamma}_{1}\right)$ (4.4)
[…]
@@
Consider nuisance functions
$\pi_{1}(X)$ $=\frac{1}{p(A=0\mid X)},\quad\pi_{2}(M,X)=\frac{p(M\mid A=0,X)}{p(M,A=1\mid X)},$
and recall that
$\mu_{1}(X,M)=\mathbb{E}[Y\mid X,M,A=1],\quad\mu_{10}(X)=\mathbb{E}[\mu_{1}(X,M)\mid X,A=0].$
In a recent work, Liu and Ghassami, (2024) showed that writing the the estimator in Equation (4.2), as
$\hat{\theta}^{mr}=\mathbb{E}_{n}\left[\hat{\mu}_{10}(X)+(1-A)\hat{\pi}_{1}(X)\left\{\hat{\mu}_{1}(M,X)-\hat{\mu}_{10}(X)\right\}+A\hat{\pi}_{2}(M,X)\left\{Y-\hat{\mu}_{1}(M,X)\right\}\right],$
the bias has the following form:
$\mathbb{E}\left[\hat{\theta}^{mr}\right]-\theta_{0}=$ $\mathbb{E}[\left\{(1-A)\hat{\pi}_{1}(X)-1\right\}\left\{\mu_{10}(X)-\hat{\mu}_{10}(X)\right\}$
$\qquad+\left\{A\hat{\pi}_{2}(M,X)-(1-A)\hat{\pi}_{1}(X)\right\}\left\{\mu_{1}(M,X)-\hat{\mu}_{1}(M,X)\right\}].$ (4.5)
[…]
@@
$\leq\mathbb{E}\left[\left\|X^{\top}\left\{A\hat{\pi}_{2}(M,X)-1\right\}\right\|_{\infty}\left\|\beta_{1}-\hat{\beta}_{1}\right\|_{1}\right]$
$+\mathbb{E}\left[\left\{\left\|M^{\top}\left\{A\hat{\pi}_{2}(M,X)\right\}-(1-A)\hat{\pi}_{1}(X)\right\|_{\infty}+\left\|(\mathbf{B}_{0}X)^{\top}\left\{(1-A)\hat{\pi}_{1}(X)-1\right\}\right\|_{\infty}\right\}\left\|\gamma_{1}-\hat{\gamma}_{1}\right\|_{1}\right]$
$+\mathbb{E}\left[\left\|X^{\top}\left\{(1-A)\hat{\pi}_{1}(X)-1\right\}\right\|_{\infty}\left\|\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}-\widehat{\mathbf{B}_{0}^{\top}\hat{\gamma}_{1}}\right\|_{1}\right].$
@@
$|\Delta_{n}|$ $\leq\|a-\mathbf{W}_{t}^{\top}\tau_{1}\|_{\infty}\|\hat{\phi}_{1}-\phi_{1}\|_{1}+\|(\hat{\mathbf{B}}_{0}-\mathbf{B}_{0})\bar{X}\|_{\infty}\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}+\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\|_{\infty}\|\hat{b}-b\|_{1}$
$=\|a-\mathbf{W}_{t}^{\top}\tau_{1}\|_{\infty}\|\hat{\beta}_{1}-\beta_{1}\|_{1}$
$\quad+\left(\|a-\mathbf{W}_{t}^{\top}\tau_{1}\|_{\infty}+\|(\hat{\mathbf{B}}_{0}-\mathbf{B}_{0})\bar{X}\|_{\infty}\right)\|\hat{\gamma}_{1}-\gamma_{1}\|_{1}$
@@
It is apparent that the bias structures of $\hat{\theta}^{mr}$ and $\hat{\theta}_{0}$ are similar, as both of them are the sum of three terms, and each of them is of the form of the product of errors.
Relation to Balancing: We also draw a connection between our debiasing technique and the balancing techniques in the literature of causal inference. In the influence function-based estimator for ATE, the correction term (the term added to the initial estimator to reduce bias) contains the propensity score as a nuisance function. Imai and Ratkovic, (2014) showed that the propensity score satisfies a certain balancing property, based on which they proposed a balancing estimator for the propensity score. Zubizarreta, (2015) proposed an optimization-based approach for achieving the said balance. In the high-dimensional setting, Athey et al., (2018) illustrated that the standard debiasing technique of a penalized regression-based estimator (e.g., the method proposed earlier in (Javanmard and Montanari, , 2014; Van de Geer et al., , 2014; Zhang and Zhang, , 2014)) is inherently related to the balancing approach of Zubizarreta, (2015). However, none of these approaches is directly applicable in the presence of mediators for assessing the mediation functional. Recently, Liu and Ghassami, (2024) presented a counterpart of the balancing property and balancing estimator of Imai and Ratkovic, (2014) for mediation functional in a fixed-dimensional setting. The counterpart of the terms involved in the balancing approach of (Liu and Ghassami, , 2024) in our setting are $\|\bar{X}-\mathbf{X}_{c}^{\top}\tau_{2}\|_{\infty}$ and $\|a-\mathbf{W}_{t}^{\top}\tau_{1}\|_{\infty}$. Hence, the two optimizations (3.1) and (3.2) performed in our proposed estimation strategy can be viewed as adaptations of the balancing of Liu and Ghassami, (2024) in the UHD setting.